A certain amount of gas occupies 5dm3 at 2 atm and 10degrees Celsius. Calculate the number of moles present.
[ R= 0.082]
1
Expert's answer
2020-04-24T13:36:02-0400

p x V = n x R x T

5 dm3 = 5 L

10 °C = 10 + 273 = 283 K

2 atm x 5 L = n x 0.082 atm•L/(mol•K) x 283 K

n = 2 atm x 5 L / (0.082 atm•L/(mol•K) x 283 K) = 0.43 mol